OpenStudy (anonymous):

Maya puts $4000 into a bank account that is earning 3.5% interest. How much will he have after ten years if the interest is compounded monthly? daily? continuously?

OpenStudy (anonymous):

\[4000(1+\frac{.035}{12})^{12\times 10}\] for the first one

OpenStudy (anonymous):

\[4000(1+\frac{.035}{365})^{365\times 10}\] for the second

OpenStudy (anonymous):

\[4000e^{.035\times 10}\] for the last one calculators for all

OpenStudy (anonymous):

Geometric series